on january 1 2011, Maya co issued 60,000 additional shares. This brought the total number of shares in issue to 660,000. In the year to 30 April 2011, the profit before tax was 847,400 and the tax charge was 47,800. what was earning per share for the year to 30 april 2011
1
Expert's answer
2016-02-10T00:01:18-0500
If the total number of shares in issue is 660,000, the profit before tax was 847,400 and the tax charge was 47,800, then earning per share for the year is:
Earning per share = (Profit before tax - Tax)/Number of shares = (847,400 - 47,800)/660,000 = 799,600/660,000 = $1.21 per share.